Q119 / 160 · WBJEE 2015 · Mathematics
Suppose a machine produces metal parts that contain some defective parts witl probability $0.05 .$ How many parts should be produced in order that the probability 0 atleast one part being defective is $1 / 20$ more? (Given that, $\log _{10} 95=1.977$ an $\log _{10} 2=0.3$ )
- A11
- B12
- Correct answerC15
- Correct answerD14
Solution
Given probability of delective pat $=0.05=\frac{1}{20}$
Probability of non-defective part $=1-0.05=0.95=\frac{19}{20}$
We know that, $P(X=r)={ }^{n} Cr p^{r} q^{n-r}$
where, $\quad p=\frac{1}{20}, q=\frac{19}{20}$
$r \geq 1$ and $n=?$
Also, $P(X \geq 1) \geq \frac{1}{2}$
$\Rightarrow 1-P(X=0) \geq \frac{1}{2}$
$\Rightarrow 1-{ }^{n} C_{0}\left(\frac{1}{20}\right)^{0}\left(\frac{19}{20}\right)^{n-0} \geq \frac{1}{2}$
$\Rightarrow \quad 1-\frac{1}{2} \geq\left(\frac{19}{20}\right)^{n}$
$\Rightarrow$
$\frac{1}{2} \geq\left(\frac{19}{20}\right)^{n}$
$\Rightarrow$
$\frac{1}{2} \geq\left(\frac{95}{100}\right)^{n}$
$\Rightarrow \quad \log 2^{-1} \geq n[\log 95-\log 100 \mid$
$\Rightarrow \quad-\log 2 \geq n[\log 95-2]$
$\Rightarrow \quad-(0.3) \geq n[1977-2]$
$\Rightarrow \quad n \geq \frac{03}{0.023}$
$\Rightarrow \quad n \geq \frac{300}{23}$
$\Rightarrow \quad n \geq 1304$
$\therefore \quad n=14,15$
Hence, the correct answer is options: 'C', 'D'.